OpenStudy (anonymous):

a,b,x are Natural numbers. a = b then a+x=b+x. can this be proved in some way ?

OpenStudy (anonymous):

begin with a = b let x = 5 add 5 to both sides a + 5 = b + 5 replace 5 with x a + x = b + x

OpenStudy (anonymous):

yes, if a = b, then obvioulsy the second equation is true. just plug in numbers like: 2=2; therefore, 2+x=2+x!

OpenStudy (anonymous):

I would like a proof. Not an example

OpenStudy (mathmate):

hint: proof by contradiction.

OpenStudy (anonymous):

begin with a = b let x = some natural number add x to both sides a + x = b + x

OpenStudy (anonymous):

What are the axioms from which i can prove this ?